ROTATIONALLY RESOLVED ABSORPTION OF O_{2} IN THE VISIBLE AT 90K

\underline{YASNAHIR PEREZ-DELGADO} and CARLOS E. MANZANARES, Department of Chemistry and Biochemistry, Baylor University, Waco, Texas, 76798.

The atmospheric red bands of molecular oxygen due to b^{1} ∑_{g}^{+}←X^{3} ∑_{g}^{−} magnetic dipole electronic transition have been obtained using the phase shift and pulsed cavity ring down technique. A low temperature cell was designed and adapted to the CRD experiment to measure absorption bands at any temperature between 10 K and 298 K using liquid He or liquid N_{2} as cryogens. The spectra were obtained at room temperature and at 90 K. The rotationally resolved spectra of the (0,0) vibration band, A-band, at 762nm and the (2,0) vibration band, γ-band, at 628nm were obtained and compared with the simulated spectrum. This information was used to confirm the temperature of the cell.